Each matrix in an equation of the form AX=B has a name
Coefficient matrix A Vaiable Matriz X Constant Matrix B
2 1 -1 x 5
3 -1 2 * y = -1
1 -1 -1 z 19
Find the product matrix
2 1 -1 x
3 -1 2 y
1 -1 -1 z

Answer by venugopalramana(3286)

Each matrix in an equation of the form AX=B has a name
Coefficient matrix A Vaiable Matriz X Constant Matrix B
2 1 -1 x 5
3 -1 2 * y = -1
1 -1 -1 z 19
IT MEANS ASSUMING SOME POSSIBLE TYPOS IN YOUR PROBLEM
2X+1Y-1Z = 5..............................I
3X-1Y+2Z = -1..............................II
1X-1Y-1Z = 19...................................III

Find the product matrix
2 1 -1 x
3 -1 2 y
1 -1 -1 z
2X+1Y-1Z.............
3X-1Y+2Z..............
1Z-1Y-1Z...................